# Model-independent predictions for decays of double-heavy hadrons into pairs of heavy hadrons

### v2

@@ III.1 Coupling Potentials
In the B-O approximation, a double-heavy hadron state can be expanded in terms of light-QCD states in the presence of static sources at the positions $+\frac{1}{2}\bm{r}$ and $-\frac{1}{2}\bm{r}$. The light-QCD state can be chosen to have light-QCD angular-momentum/parity $j^{\pi}$ and B-O quantum numbers $\lambda$ and $\eta$. We denote the state by a ket $\lvert(+,-)j^{\pi},\lambda,\eta\rangle$ in the text and by a ket $\Bigl{|}(+,-)\begin{smallmatrix}j^{\pi}\\[1.99997pt] \lambda\end{smallmatrix},\eta\Bigr{\rangle}$ in equations. The multiplet corresponding to the $12j+$ 1 values of $\lambda$ will be labeled more concisely as $(+,-)j^{\pi}$. The double-heavy hadron state can be expressed in terms of integrals over $\bm{r}$ of these $\bm{r}$-dependent kets multiplied by wavefunctions.
The decay of a double-heavy hadron into a pair of heavy hadrons can proceed through the transition of a light-QCD state of the double-heavy hadron into a light-QCD state of a static-hadron pair with the same B-O quantum numbers. Up to an appropriate normalization coefficient, the B-O transition amplitude as a function of $\bm{r}$ can be defined as the time derivative of the transition amplitude between the light-QCD state $\lvert(+,-)j^{\pi},\lambda,\eta\rangle$ of the double-heavy hadron and the light-QCD state $\bigl{\lvert}\bigl{(}j_{1}^{\pi_{1}}(+),j_{2}^{\pi_{2}}(-)\bigr{)}j^{\prime},\lambda,\eta\bigr{\rangle}$ of a pair of static hadrons,
@@
with $\mathcal{T}_{t,t_{0}}$ the evolution operator from time $t_{0}$ to time $t$. The bra and the ket on the right side are functions of $\bm{r}$. The rotational symmetry of QCD implies that the transition amplitude in Eq. (25) is a function of the distance $r=\lvert\bm{r}\rvert$ only. The subscripts on $g$ are quantum numbers $\lambda$ and $\eta$ that are conserved in light QCD with two static color sources.
The B-O transition amplitude in Eq. (25) induces decays of a double-heavy hadron into pairs of heavy hadrons. The decay widths can be calculated by solving a Schrödinger equation with coupled double-heavy-hadron and heavy-hadron-pair channels. The total angular momentum of the double-heavy system is
@@
with $\bm{L}_{Q}$ the orbital-angular-momentum vector of $Q$ and $Q$ or $Q$ and $\bar{Q}$. As shown in Ref. [21] using the diabatic B-O framework, including all the appropriate coupled channels automatically ensures conservation of total angular momentum and parity. The heavy (anti)quark spins are conserved up to corrections suppressed by $1/m_{Q}$. Since the total-heavy-spin vector $\bm{S}_{Q}$ is conserved in the heavy-quark limit, the Born-Oppenheimer angular-momentum vector
@@
is also conserved in the heavy-quark limit. We denote the quantum numbers for the B-O angular-momentum vector $\bm{L}$ by $(L,m_{L})$. Double-heavy hadrons form approximately degenerate B-O multiplets labeled by the angular-momentum/parity $L^{P}$ and other quantum numbers. We denote the quantum number for the orbital-angular-momentum vector $\bm{L}_{Q}$ by $L_{Q}$. The orbital angular momentum is in general not conserved in the heavy-quark limit.
The hadron-pair states on both sides of Eq. (24) can be multiplied by orbital-angular-momentum states $\lvert L_{Q}^{\prime},m^{\prime}\rangle$ and then projected onto total-angular-momentum states $\lvert J,m_{J}\rangle$. The states on the right side of the equation can then be decomposed into states with B-O angular momentum $L$ using the angular-momentum recoupling identity
@@
where $\left\{\begin{smallmatrix}j_{1}&j_{2}&j_{3}\\ j_{4}&j_{5}&j_{6}\end{smallmatrix}\right\}$ is a Wigner 6-$j$ symbol. The states can then be projected onto states with definite $[\mathcal{C}]\mathcal{P}$-parity $\eta^{\prime}$:
|$\biggl{|}\Bigl{\{}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1},\Bigl{(}\tfrac{1}{2}^{\pm},j_{2}^{\pi_{2}}\Bigr{)}J_{2}\Bigr{]}S,L_{Q}^{\prime}\Bigr{\}}$ $\begin{matrix}J\\[-5.0pt] m_{J}\end{matrix},\eta^{\prime}\biggr{\rangle}=N(-1)^{2j_{1}+L_{Q}^{\prime}+J}\sqrt{\tilde{J}_{1}\tilde{J}_{2}\tilde{S}}$ (31)
×$\times\sum_{j^{\prime},S_{Q},L}(-1)^{j^{\prime}+S_{Q}}\sqrt{\tilde{\jmath}^{\prime}\tilde{S}_{Q}\tilde{L}}\begin{Bmatrix}S_{Q}&j^{\prime}&S\\ L_{Q}^{\prime}&J&L\\ \end{Bmatrix}\begin{Bmatrix}\frac{1}{2}&\frac{1}{2}&S_{Q}\\ j_{1}&j_{2}&j^{\prime}\\ J_{1}&J_{2}&S\\ \end{Bmatrix}$ (37)
×|$\times\biggl{|}\Bigl{\{}\Bigl{(}\tfrac{1}{2}^{+},\tfrac{1}{2}^{\pm}\Bigr{)}S_{Q},\Bigl{[}\Bigl{(}j_{1}^{\pi_{1}},j_{2}^{\pi_{2}}\Bigr{)}j^{\prime},L_{Q}^{\prime}\Bigr{]}L\Bigr{\}}\begin{matrix}J\\[-5.0pt] m_{J}\end{matrix},\eta^{\prime}\biggr{\rangle},$ (40)
where the normalization coefficient $N$ is 1 unless the two static hadrons $j_{1}^{\pi_{1}}$ and $j_{2}^{\pi_{2}}$ are identical (for $(\bm{3},\bm{3})$ sources) or charge conjugates (for $(\bm{3},\bm{3}^{\ast})$ sources). Even in this case, $N$ differs from 1 only if $J_{1}\neq J_{2}$, in which case $N=\sqrt{2}$. This dependence of $N$ on the quantum numbers has been suppressed in Eq. (40).
@@
with $L_{Q}$ and $L_{Q}^{\prime}$ the orbital angular momenta for the double-heavy hadron and the heavy-hadron pair, respectively. The superscript on $G$ is the quantum number $P$ that is exactly conserved. The subscripts on $G$ are quantum numbers $L$ and $\eta$ that are conserved in the heavy-quark limit. Note that the quantum numbers $P$, $\pi$, and $L_{Q}$ are constrained according to
$P=\pm\pi(-1)^{L_{Q}}.$ (42)
@@
The right side is the product of the parity quantum numbers for the $QQ$ or $Q\bar{Q}$, the light-QCD state, and the orbital-angular-momentum state.
@@
In contrast to the coupling potentials in Ref. [21], the transition amplitudes in Eq. (41) do not take into account the heavy (anti)quark spins but they have the advantage that they allow the conservation of $L$ to be exploited. The corresponding transition operator can be obtained by multiplying Eq. (41) on the left by the ket $\bigl{\lvert}\bigl{[}\bigl{(}j_{1}^{\pi_{1}},j_{2}^{\pi_{2}}\bigr{)}j^{\prime},L_{Q}^{\prime}\bigr{]}L,m_{L},\eta\bigr{\rangle}$ and on the right by the bra $\bigl{\langle}\bigl{(}j^{\pi},L_{Q}\bigr{)}L,m_{L},\eta\bigr{\rvert}$ and then summing over $j^{\prime}$, $L_{Q}^{\prime}$, $L_{Q}$, $L$, $m_{L}$, and $\eta$. Note that we have suppressed the arguments (+)(+)( + ) and ()(-)( - ) of $\bigl{(}j_{1}^{\pi_{1}},j_{2}^{\pi_{2}}\bigr{)}j^{\prime}$ and the argument (+,)(+,-)( + , - ) of $j^{\pi}$ since we have projected the state of a $QQ$ or $Q\bar{Q}$ pair at the positions $+\frac{1}{2}\bm{r}$, $-\frac{1}{2}\bm{r}$ onto states with orbital angular momenta $L_{Q}^{\prime}$ and $L_{Q}$, respectively, and distance $r$.
@@
The transition operator with the heavy (anti)quark spins included can be obtained by multiplying the transition operator corresponding to the transition amplitudes in Eq. (41) by the identity operator for the heavy (anti)quark spins, which can be expressed as a sum of projectors $\lvert S_{Q},m\rangle\langle S_{Q},m\rvert$ over the quantum numbers $S_{Q}$ and $m$. The coupling potentials are obtained by taking the matrix element between the bra corresponding to the heavy-hadron-pair state in Eq. (40) and the double-heavy-hadron state $\Bigl{\lvert}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},\tfrac{1}{2}^{\pm}\Bigr{)}S_{Q},\Bigl{(}j^{\pi},L_{Q}\Bigr{)}L\Bigr{]}J,m_{J},\eta^{\prime}\Bigr{\rangle}$. Note that the state in Eq. (40) contains a sum over the quantum numbers $j^{\prime}$, $S_{Q}$, and $L$. The matrix element simplifies in the heavy-quark limit, since $S_{Q}$ and $L$ are conserved in this limit. Thus only the sum over $j^{\prime}$ survives in the heavy-quark limit of the matrix element. The resulting coupling potentials reduce to
@@
$V^{J,P,\eta^{\prime}}_{S_{Q},L,\eta}\Bigl{(}j^{\pi},L_{Q}\to\bigl{[}\bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\bigr{)}J_{1},\bigl{(}\tfrac{1}{2}^{\pm},j_{2}^{\pi_{2}}\bigr{)}J_{2}\bigr{]}S,L_{Q}^{\prime}\Bigr{)}=N(-1)^{2j_{1}+S_{Q}+L_{Q}^{\prime}+J}\sqrt{\tilde{J}_{1}\tilde{J}_{2}\tilde{S}\tilde{S}_{Q}\tilde{L}}\\ \times\sum_{j^{\prime}}(-1)^{j^{\prime}}\sqrt{\tilde{\jmath}^{\prime}}\begin{Bmatrix}S_{Q}&j^{\prime}&S\\ L_{Q}^{\prime}&J&L\\ \end{Bmatrix}\begin{Bmatrix}\frac{1}{2}&\frac{1}{2}&S_{Q}\\ j_{1}&j_{2}&j^{\prime}\\ J_{1}&J_{2}&S\\ \end{Bmatrix}G_{L,\eta}^{P}\bigl{(}j^{\pi},L_{Q}\to(j_{1}^{\pi_{1}},j_{2}^{\pi_{2}})j^{\prime},L_{Q}^{\prime}\bigr{)}.$ (43)
@@
$\eta^{\prime}=\pm\eta(-1)^{S_{Q}}.$ (44)
The right side is the product of the $[\mathcal{C}]\mathcal{P}$ quantum numbers for the $QQ$ or $Q\bar{Q}$, the light-QCD state, and the total-heavy-spin state.
@@ III.2 Selection Rules
The transition of a light-QCD state $(+,-)j^{\pi}$ of a double-heavy hadron into the state of a pair of static hadrons $j_{1}^{\pi_{1}}(+)$ and $j_{2}^{\pi_{2}}(-)$ is forbidden if the B-O quantum numbers of the former, $\lambda$ and $\eta$ (and $\epsilon$), cannot be obtained from any of the static-hadron-pair states $(j_{1}^{\pi_{1}}(+),j_{2}^{\pi_{2}}(-))j^{\prime}$ with $j^{\prime}$ ranging from $\lvert j_{1}-j_{2}\rvert$ to $j_{1}+j_{2}$. This implies model-independent Born-Oppenheimer selection rules.
@@ III.2.1 Selection Rules for $\lambda$
If the two static hadrons are identical (for $(\bm{3},\bm{3})$ sources) or charge conjugates (for $(\bm{3},\bm{3}^{\ast})$ sources), this B-O selection rule becomes more restrictive upon taking into account conservation of $\eta$ in accordance with Eq. (8). In this case, which requires $j_{2}=j_{1}$, the range of $j^{\prime}$ is from 0 to $2j_{1}$. For $(\bm{3},\bm{3})$ sources, the largest value $j^{\prime}=2j_{1}$ is possible only if $P=(-1)^{2j_{1}}$. For $(\bm{3},\bm{3}^{\ast})$ sources, the largest value $j^{\prime}=2j_{1}$ is possible only if $CP=+1$ since $4j_{1}$ is always even. For the opposite value of $\eta$, the largest value of $j^{\prime}$ is $12j_{1}-$ 1. If $P=-(-1)^{2j_{1}}$ (for $(\bm{3},\bm{3})$ sources) or if $CP=-1$ for ($(\bm{3},\bm{3}^{\ast})$ sources), the selection rule for $\lambda$ in Eq. (45) is replaced by [^4] Note that if $j_{1}=j_{2}=0$, Eq. (45) becomes $\lambda=0$ and Eq. (46) does not apply.
@@
III.2.2 Selection Rules for $\Sigma_{\eta}^{\epsilon}$
In the special case $\lambda=0$, the B-O quantum numbers are $\Sigma_{\eta}^{\epsilon}$ and there is a B-O selection rule associated with conservation of $\epsilon$. Following Eq. (13), the static-hadron-pair states with integer $j^{\prime}$ and $\lambda=0$ are eigenstates of reflections $\mathcal{R}$ with eigenvalue $\pi_{1}\pi_{2}(-1)^{j^{\prime}}$. Thus, if the light-QCD state of the double-heavy hadron is $\Sigma_{\eta}^{\epsilon}$, $\epsilon$ must satisfy

### v3

@@ III.1 Coupling Potentials
In the B-O approximation, a hidden-heavy hadron state can be expanded in terms of light-QCD states in the presence of static $\bm{3}$ and $\bm{3}^{\ast}$ sources at the positions $+\frac{1}{2}\bm{r}$ and $-\frac{1}{2}\bm{r}$. The light-QCD state can be chosen to have B-O quantum numbers $\lambda$ and $\eta$. In the $\bm{r}\to 0$ limit, the light-QCD state is dominated by light-QCD angular-momentum/parity $j^{\pi}$. We denote the state by a ket $\lvert(+,-)j^{\pi},\lambda,\eta\rangle$ in the text and by a ket $\Bigl{|}(+,-)\begin{smallmatrix}j^{\pi}\\[1.99997pt] \lambda\end{smallmatrix},\eta\Bigr{\rangle}$ in equations. The multiplet corresponding to the $12j+$ 1 values of $\lambda$ will be labeled more concisely as $(+,-)j^{\pi}$. The hidden-heavy hadron state can be expressed in terms of integrals over $\bm{r}$ of these $\bm{r}$-dependent kets multiplied by wavefunctions.
The decay of a hidden-heavy hadron into a pair of heavy hadrons can proceed through the transition of a light-QCD state of the hidden-heavy hadron into a light-QCD state of a static-hadron pair with the same B-O quantum numbers. Up to an appropriate normalization coefficient, the B-O transition amplitude as a function of $\bm{r}$ can be defined as the time derivative of the transition amplitude between the light-QCD state $\lvert(+,-)j^{\pi},\lambda,\eta\rangle$ of the hidden-heavy hadron and the light-QCD state $\bigl{\lvert}\bigl{(}j_{1}^{\pi_{1}}(+),j_{2}^{\pi_{2}}(-)\bigr{)}j^{\prime},\lambda,\eta\bigr{\rangle}$ of a pair of static hadrons,
@@
with $\mathcal{T}_{t,t_{0}}$ the evolution operator from time $t_{0}$ to time $t$. The bra and the ket on the right side are functions of $\bm{r}$. The rotational symmetry of QCD implies that the transition amplitude in Eq. (25) is a function of the distance $r=\lvert\bm{r}\rvert$ only. The subscripts on $g$ are quantum numbers $\lambda$ and $\eta$ that are conserved in light QCD with static $\bm{3}$ and $\bm{3}^{\ast}$ sources.
The B-O transition amplitude in Eq. (25) induces decays of a hidden-heavy hadron into pairs of heavy hadrons. The decay widths can be calculated by solving a Schrödinger equation with coupled hidden-heavy-hadron and heavy-hadron-pair channels. The total angular momentum of the hidden-heavy system is
@@
with $\bm{L}_{Q}$ the orbital-angular-momentum vector of $Q$ and $\bar{Q}$. As shown in Ref. [21] using the diabatic B-O framework, including all the appropriate coupled channels automatically ensures conservation of total angular momentum and parity. The heavy (anti)quark spins are conserved up to corrections suppressed by $1/m_{Q}$. Since the total-heavy-spin vector $\bm{S}_{Q}$ is conserved in the heavy-quark limit, the Born-Oppenheimer angular-momentum vector
@@
is also conserved in the heavy-quark limit. We denote the quantum numbers for the B-O angular-momentum vector $\bm{L}$ by $(L,m_{L})$. Hidden-heavy hadrons form approximately degenerate B-O multiplets labeled by the angular-momentum/parity $L^{P}$ and other quantum numbers. We denote the quantum number for the orbital-angular-momentum vector $\bm{L}_{Q}$ by $L_{Q}$. In general, the orbital angular momentum is not conserved in the heavy-quark limit.
The hadron-pair states on both sides of Eq. (24) depend on the vector $\bm{r}$. They can be expanded in terms of orbital-angular-momentum states $\lvert L_{Q}^{\prime},m^{\prime}\rangle$ and then projected onto total-angular-momentum states $\lvert J,m_{J}\rangle$. Furthermore, the states on the right side of the equation can be decomposed into states with B-O angular momentum $L$ using the angular-momentum recoupling identity
@@
where $\left\{\begin{smallmatrix}j_{1}&j_{2}&j_{3}\\ j_{4}&j_{5}&j_{6}\end{smallmatrix}\right\}$ is a Wigner 6-$j$ symbol. Finally, the states can be projected onto states with definite $\mathcal{C}\mathcal{P}$-parity $\eta^{\prime}$:
|$\biggl{|}\Bigl{\{}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\Bigr{)}J_{1},\Bigl{(}\tfrac{1}{2}^{-},j_{2}^{\pi_{2}}\Bigr{)}J_{2}\Bigr{]}S,L_{Q}^{\prime}\Bigr{\}}$ $\begin{matrix}J\\[-5.0pt] m_{J}\end{matrix},\eta^{\prime}\biggr{\rangle}=N(-1)^{2j_{1}+L_{Q}^{\prime}+J}\sqrt{\tilde{J}_{1}\tilde{J}_{2}\tilde{S}}$
×$\times\sum_{j^{\prime},S_{Q},L}(-1)^{j^{\prime}+S_{Q}}\sqrt{\tilde{\jmath}^{\prime}\tilde{S}_{Q}\tilde{L}}\begin{Bmatrix}S_{Q}&j^{\prime}&S\\ L_{Q}^{\prime}&J&L\\ \end{Bmatrix}\begin{Bmatrix}\frac{1}{2}&\frac{1}{2}&S_{Q}\\ j_{1}&j_{2}&j^{\prime}\\ J_{1}&J_{2}&S\\ \end{Bmatrix}$
×|$\times\biggl{|}\Bigl{\{}\Bigl{(}\tfrac{1}{2}^{+},\tfrac{1}{2}^{-}\Bigr{)}S_{Q},\Bigl{[}\Bigl{(}j_{1}^{\pi_{1}},j_{2}^{\pi_{2}}\Bigr{)}j^{\prime},L_{Q}^{\prime}\Bigr{]}L\Bigr{\}}\begin{matrix}J\\[-5.0pt] m_{J}\end{matrix},\eta^{\prime}\biggr{\rangle},$ (29)
where the normalization coefficient $N$ is 1 unless the two static hadrons $j_{1}^{\pi_{1}}$ and $j_{2}^{\pi_{2}}$ are charge conjugates. Even in this case, $N$ differs from 1 only if $J_{1}\neq J_{2}$, in which case $N=\sqrt{2}$. This dependence of $N$ on the quantum numbers has been suppressed in Eq. (29). Note that we have dropped the arguments (+)(+)( + ) and ()(-)( - ) since we have expanded the state of two heavy hadrons at the positions $+\frac{1}{2}\bm{r}$ and $-\frac{1}{2}\bm{r}$ into states with orbital angular momentum $L_{Q}^{\prime}$ and distance $r$.
@@
with $L_{Q}$ and $L_{Q}^{\prime}$ the orbital angular momenta for the hidden-heavy hadron and the heavy-hadron pair, respectively. The superscript on $G$ is the quantum number $P$ that is exactly conserved. The subscripts on $G$ are quantum numbers $L$ and $\eta$ that are conserved in the heavy-quark limit. Note that the quantum numbers $P$, $\pi$, and $L_{Q}$ are constrained according to
$P=\pi(-1)^{L_{Q}+1}.$ (31)
@@
The right side is the product of the parity quantum numbers for the light-QCD state and the $Q\bar{Q}$ total-angular-momentum state.
@@
In contrast to the coupling potentials in Ref. [21], the transition amplitudes in Eq. (30) do not take into account the heavy (anti)quark spins but they have the advantage that they allow the conservation of $L$ to be exploited. The corresponding transition operator can be obtained by multiplying Eq. (30) on the left by the ket $\bigl{\lvert}\bigl{[}\bigl{(}j_{1}^{\pi_{1}},j_{2}^{\pi_{2}}\bigr{)}j^{\prime},L_{Q}^{\prime}\bigr{]}L,m_{L},\eta\bigr{\rangle}$ and on the right by the bra $\bigl{\langle}\bigl{(}j^{\pi},L_{Q}\bigr{)}L,m_{L},\eta\bigr{\rvert}$ and then summing over $j^{\prime}$, $L_{Q}^{\prime}$, $L_{Q}$, $L$, $m_{L}$, and $\eta$. Note that we have dropped the argument (+,)(+,-)( + , - ) of $j^{\pi}$ since we have expanded the state of $Q$ and $\bar{Q}$ at the positions $+\frac{1}{2}\bm{r}$ and $-\frac{1}{2}\bm{r}$ into states with orbital angular momentum $L_{Q}$ and distance $r$.
@@
The transition operator with the heavy (anti)quark spins included can be obtained by multiplying the transition operator corresponding to the transition amplitudes in Eq. (30) by the identity operator for the heavy (anti)quark spins, which can be expressed as a sum of projectors $\lvert S_{Q},m\rangle\langle S_{Q},m\rvert$ over the quantum numbers $S_{Q}$ and $m$. The coupling potentials are obtained by taking the matrix element between the bra corresponding to the heavy-hadron-pair state in Eq. (29) and the hidden-heavy-hadron state $\Bigl{\lvert}\Bigl{[}\Bigl{(}\tfrac{1}{2}^{+},\tfrac{1}{2}^{-}\Bigr{)}S_{Q},\Bigl{(}j^{\pi},L_{Q}\Bigr{)}L\Bigr{]}J,m_{J},\eta^{\prime}\Bigr{\rangle}$. Note that the state in Eq. (29) contains a sum over the quantum numbers $j^{\prime}$, $S_{Q}$, and $L$. The matrix element simplifies in the heavy-quark limit, since $S_{Q}$ and $L$ are conserved in this limit. Thus only the sum over $j^{\prime}$ survives in the heavy-quark limit of the matrix element. The resulting coupling potentials reduce to
@@
$V^{J,P,\eta^{\prime}}_{S_{Q},L,\eta}\Bigl{(}j^{\pi},L_{Q}\to\bigl{[}\bigl{(}\tfrac{1}{2}^{+},j_{1}^{\pi_{1}}\bigr{)}J_{1},\bigl{(}\tfrac{1}{2}^{-},j_{2}^{\pi_{2}}\bigr{)}J_{2}\bigr{]}S,L_{Q}^{\prime}\Bigr{)}=N(-1)^{2j_{1}+S_{Q}+L_{Q}^{\prime}+J}\sqrt{\tilde{J}_{1}\tilde{J}_{2}\tilde{S}\tilde{S}_{Q}\tilde{L}}\\ \times\sum_{j^{\prime}}(-1)^{j^{\prime}}\sqrt{\tilde{\jmath}^{\prime}}\begin{Bmatrix}S_{Q}&j^{\prime}&S\\ L_{Q}^{\prime}&J&L\\ \end{Bmatrix}\begin{Bmatrix}\frac{1}{2}&\frac{1}{2}&S_{Q}\\ j_{1}&j_{2}&j^{\prime}\\ J_{1}&J_{2}&S\\ \end{Bmatrix}G_{L,\eta}^{P}\bigl{(}j^{\pi},L_{Q}\to(j_{1}^{\pi_{1}},j_{2}^{\pi_{2}})j^{\prime},L_{Q}^{\prime}\bigr{)}.$ (32)
@@
$\eta^{\prime}=\eta(-1)^{S_{Q}+1}.$ (33)
The right side is the product of the $\mathcal{C}\mathcal{P}$ quantum numbers for the light-QCD state and the $Q\bar{Q}$ total-heavy-spin state.
@@ III.2 Selection Rules
The transition of the light-QCD state $(+,-)j^{\pi}$ of a hidden-heavy hadron into the state of a pair of static hadrons $j_{1}^{\pi_{1}}(+)$ and $j_{2}^{\pi_{2}}(-)$ is forbidden if the B-O quantum numbers of the former, $\lambda$ and $\eta$ (and $\epsilon$), cannot be obtained from any of the static-hadron-pair states $(j_{1}^{\pi_{1}},j_{2}^{\pi_{2}})j^{\prime}$ with $j^{\prime}$ ranging from $\lvert j_{1}-j_{2}\rvert$ to $j_{1}+j_{2}$. This implies model-independent Born-Oppenheimer selection rules.
@@ III.2.1 Selection Rules for $\lambda$
If the two static hadrons are charge conjugates, this B-O selection rule becomes more restrictive upon taking into account conservation of $\eta$ in accordance with Eq. (8). In this case, which requires $j_{2}=j_{1}$, the range of $j^{\prime}$ is from 0 to $2j_{1}$. The largest value $j^{\prime}=2j_{1}$ is possible only if $\eta=+1$ since $4j_{1}$ is always even. If $\eta=-1$, the largest value of $j^{\prime}$ is $12j_{1}-$ 1 and the selection rule for $\lambda$ in Eq. (34) is replaced by [^3] Note that if $j_{1}=j_{2}=0$, Eq. (34) becomes $\lambda=0$ and Eq. (35) does not apply.
@@
III.2.2 Selection Rules for $\epsilon$
In the special case $\lambda=0$, the B-O quantum numbers are $\Sigma_{\eta}^{\epsilon}$ and there is a B-O selection rule associated with conservation of $\epsilon$. Following Eq. (13), the static-hadron-pair states with integer $j^{\prime}$ and $\lambda=0$ are eigenstates of reflections $\mathcal{R}$ with eigenvalue $\pi_{1}\pi_{2}(-1)^{j^{\prime}}$. Thus, if the light-QCD state of the hidden-heavy hadron is $\Sigma_{\eta}^{\epsilon}$, $\epsilon$ must satisfy